A point charge 10micro couloumb distance 5cm directly above the centre of a square of side 10cm . What is the magnitude of the electric flux through the square?

Grade:12

1 Answers

Arun
5 years ago
The square can be considered as one face of a cube of edge 10 cm with a centre where charge q is placed. According to Gauss’s theorem for a cube, total electric flux is through all its six faces.
 
Hence, electric flux through one face of the cube i.e., through the square,
 
 = (1/6)(q/ epsilon)
 

Where, ∈0 = Permittivity of free space

= 8.854 × 10−12 N−1C2 m−2 q = 10 μC = 10 × 10−6 C

 

= 1.88 × 105 N m2 C−1

Therefore, electric flux through the square is 1.88 × 105 N m2C−1.